# Trans-Glasso: A Transfer Learning Approach to Precision Matrix Estimation

### v1

@@ 3.2 Refinement via Differential Network Estimation
To further enhance accuracy, we refine these initial estimators by estimating the differential networks $\Psi^{(k)}=\Omega^{(k)}-\Omega^{(0)}$, which capture the structural differences between each source and the target. This refinement step corrects for potential biases in the initial estimates.
Estimating $\Psi^{(k)}$ has been extensively studied in the differential network estimation literature [Zhao et al., 2014, Yuan et al., 2017], with a variety of good estimators. For instance, Yuan et al. [2017] proposed estimation by solving
@@
where
$L_{D}\left(\Psi;\widehat{\Sigma}^{(0)},\widehat{\Sigma}^{(k)}\right)=\frac{1}{4}\left(\left\langle\widehat{\Sigma}^{(0)}\Psi,\Psi\widehat{\Sigma}^{(k)}\right\rangle+\left\langle\widehat{\Sigma}^{(k)}\Psi,\Psi\widehat{\Sigma}^{(0)}\right\rangle\right)-\left\langle\Psi,\widehat{\Sigma}^{(0)}-\widehat{\Sigma}^{(k)}\right\rangle,$
and $\lambda^{(k)}_{\Psi}$ is a tuning parameter.
For our purpose, any reasonable differential network estimator can correct for the bias. Thus, we treat differential network estimation as a black-box algorithm, and we obtain the estimates $\widehat{\Psi}^{(k)}$.
@@
where $\widehat{\Psi}^{(0)}=0$ by definition. The final estimator Trans-Glasso (Transfer learning Graphical lasso) integrates both shared information and source-specific refinements, yielding a transfer learning approach that leverages structural similarities across datasets for improved precision matrix estimation.
Compared to Li et al. [2023a], our method is more sample efficient as it does not require sample splitting between the steps.
@@ 4.1 Optimization Algorithms
To implement Trans-Glasso, we first solve the Trans-MT-Glasso objective from Equation (3) to obtain initial estimates for $\Omega^{\star}$ and $\Gamma^{(k)\star}$. This is a constrained optimization problem that can be efficiently solved using the Alternating Direction Method of Multipliers (ADMM) [Boyd et al., 2011]. In this section, we slightly abuse notation by using superscripts to denote the iteration round and subscripts to represent the population.
@@
$X=\begin{bmatrix}\Omega_{0}\\ \Omega_{1}\\ \vdots\\ \Omega_{K}\end{bmatrix}\in\mathbb{R}^{(K+1)d\times d},\quad Y=\begin{bmatrix}\Omega\\ \Gamma_{0}\\ \Gamma_{1}\\ \vdots\\ \Gamma_{K}\end{bmatrix}\in\mathbb{R}^{(K+2)d\times d},$
and
$B=\begin{bmatrix}I_{d}&I_{d}&0&\cdots&0\\ I_{d}&0&I_{d}&\cdots&0\\ \vdots&\vdots&\vdots&&\vdots\\ I_{d}&0&0&\cdots&I_{d}\end{bmatrix}\in\mathbb{R}^{(K+1)d\times(K+2)d}.$
With this notation, the objective in Equation (3) can be reformulated as
$\text{\text{minimize}}\,\,f(X)+g(Y),\qquad\text{subject to}\,\,X=BY,$
@@
where
$Z=\begin{bmatrix}Z_{0}\\ Z_{1}\\ \vdots\\ Z_{K}\end{bmatrix}\in\mathbb{R}^{(K+1)d\times d},$
is a dual variable and $\rho$ isa penalty parameter.
@@
$Z^{(t)}=Z^{(t-1)}+\rho\left(X^{(t)}-BY^{(t)}\right).$ (11)
Note that (11) is equivalent to $Z^{(t)}_{k}=Z^{(t-1)}_{k}+\rho\left(\Omega^{(t)}_{k}-\Omega^{(t)}-\Gamma^{(t)}_{k}\right)$ for $0\leq k\leq K$. See Appendix H for detailed steps and stopping criteria.
To refine the initial estimators, we solve the D-Trace loss objective (7) for differential network estimation. Here, we use a different proximal gradient descent algorithm [Parikh and Boyd, 2014] following Zhao et al. [2022, 2019]. For simplicity, let $L_{D}(\Psi)\coloneqq L_{D}\left(\Psi\,;\,\widehat{\Sigma}^{(0)},\widehat{\Sigma}^{(k)}\right)$ for the chosen $k$. In iteration $t$, we update $\Psi^{(t-1)}$ by solving
@@
where $\eta$ is a user-specified step size. Note that $\nabla L_{D}(\cdot)$ is Lipschitz continuous with constant $\|\widehat{\Sigma}^{(0)}\otimes\widehat{\Sigma}^{(k)}\|_{2}=\|\widehat{\Sigma}^{(0)}\|_{2}\|\widehat{\Sigma}^{(k)}\|_{2}$. Thus, for $0<\eta\leq\|\widehat{\Sigma}^{(0)}\|^{-1}_{2}\|\widehat{\Sigma}^{(k)}\|^{-1}_{2}$, the proximal gradient method converges [Beck and Teboulle, 2009]. The update in (12) has a closed-form solution:
@@
where $A^{(t-1)}=\Psi^{(t-1)}-\eta\nabla L_{D}(\Psi^{(t-1)})$ and $x_{+}=\max\{0,x\}$, $x\in{\mathbb{R}}$.
Details on the optimization algorithms, including stopping criteria and descriptions, are in Section H.
@@ 4.2 Hyperparameter Selection
following Yuan et al. [2017]. After selecting $\lambda^{(k)}_{\Psi}$ and obtaining $\widehat{\Psi}^{(k)}$ for all $k\in[K]$, $\widehat{\Omega}^{(0)}$ depends on $\lambda_{\text{M}}$. Recall that $N=\sum^{K}_{k=0}n_{k}$. We choose $\lambda_{\text{M}}$ to minimize the BIC of Trans-Glasso, defined as
@@ 4.3 Identifying the Informative Set
We obtain differential network estimations $\widehat{\Psi}^{(k)}$ for all $k\in[K]$, with the hyperparameter $\lambda^{(k)}_{\Psi}$ chosen by minimizing the BIC criterion in (14). We then rank sources according to the sparsity level of $\widehat{\Psi}^{(k)}$. Let $R_{k}$ be the rank of the source $k$. For any $1\leq k_{1},k_{2}\leq K$, $|\widehat{\Psi}^{(k_{1})}|_{0}\leq|\widehat{\Psi}^{(k_{2})}|_{0}$ implies $R_{k_{1}}\leq R_{k_{2}}$. After ranking sources, we input samples into Trans-Glasso and determine the number of sources based on the cross-validation (CV) error. For $K_{\mathrm{chosen}}=0,1,\ldots,K$, we select sources $k$ with $R_{k}\leq K_{\mathrm{chosen}}$. When $K_{\mathrm{chosen}}=0$, we obtain $\widehat{\Omega}^{(0)}$ from graphical lasso [Friedman et al., 2008] using the target data alone. We then compute the CV error of $K_{\mathrm{chosen}}$ by the following procedure:
@@
We set the estimated informative set as $\widehat{\mathcal{A}}=\{k\in[K]:R_{k}\leq K^{\star}_{\text{chosen}}\}$, where $K^{\star}_{\text{chosen}}=\arg\min_{k=0,1,\ldots,K}\text{CV}(k)$. Source samples from $\widehat{\mathcal{A}}$ are used to estimate $\Omega^{(0)}$. If $\widehat{\mathcal{A}}=\emptyset$, we obtain $\widehat{\Omega}^{(0)}$ via graphical lasso on target samples. We note that sample splitting is not necessary between estimating $\mathcal{A}$ and $\Omega^{(0)}$. This procedure is called Trans-Glasso-CV. We demonstrate its empirical performance in Section 6.
@@ 5 Theoretical Analysis
In this section, we establish theoretical guarantees for the Trans-Glasso algorithm. We begin by analyzing the initial estimation step using Trans-MT-Glasso, followed by an error bound for the complete Trans-Glasso estimator. Finally, we derive a minimax lower bound, demonstrating that Trans-Glasso is minimax optimal in a wide range parameter regimes.
@@ 5.1 Analysis of Trans-MT-Glasso
The following theorem provides a high probability upper bound on the Frobenius norm error for the Trans-MT-Glasso estimator. Recall that $N=\sum^{K}_{k=0}n_{k}$.
@@
Note that the loss function defined in (4) is not strong convex with respect to the Euclidean norm. This prevents us from obtaining error guarantee for individual prevision matrix. Nevertheless, we make a key observation that the loss function exhibits strong convexity with respect to the weighted norm used in (20). See Appendix C for a detailed proof.
When we choose $\lambda_{M}\asymp\sqrt{(\log d)/N}$, the rate shown in Theorem 1 consists of two parts. The first part, which is of the order $(s\log d)/{N}$, refers to the estimation error of the shared component. In words, Trans-Glasso uses all the samples to estimate the shared component. The second part, of the order $(Kh\log d)/{N}$, relates to the estimation error of the individual components, i.e., on average, there are $N/K$ samples to estimate each individual ocmponent.
@@ 5.2 Analysis of Trans-Glasso
holds simultaneously for all $k=1,\ldots,K$ with probability at least $1-\delta$. We now establish a non-asymptotic error bound for this estimator, which combines the initial estimation error with the error from differential network estimation.
Theorem 2.
Let $\widehat{\Omega}^{(0)}$ be the Trans-Glasso estimator obtained in Equation (8). Under the same conditions as in Theorem 1, and assuming Equation (21) holds for the differential network estimators. Then with probability at least $1-2\delta$, one has
@@ 5.2.1 A differential network estimator: D-Trace loss minimization
We characterize $g^{(k)}_{\mathrm{F}}$ in the case when the D-Trace loss estimator in (7) is used. We use the tighter D-Trace loss estimator analysis from Zhao et al. [2019, 2022], Tugnait [2023].
We then have the following theorem providing a high-probability error upper bound for D-Trace loss estimator. The theorem is derived directly from Theorem 7 in Appendix D.
Theorem 3.
Suppose that Assumptions 1-2 hold. Assume that $\min_{0\leq k\leq K}n_{k}\gg h^{2}\log\left(2(K+1)d^{2}/\delta\right)$. Set
@@
for D-Trace loss estimator. Plug the above results into Theorem 2, we have the following corollary.
@@
Let $\widehat{\Omega}^{(0)}$ be obtained by Trans-Glasso (8) with the D-Trace loss estimator used in Step 1. Instate the assumptions in Theorems 1 and 3. For a given $\delta\in(0,1]$, letting
@@
holds with probability at least $1-2\delta$.
The estimation error consists of three parts: shared component estimation, individual component estimation, and differential network estimation. If $\widebar{n}\geq n_{0}$ and $M_{\Gamma}$ is bounded by a universal constant, the error scales as $\sqrt{\frac{s\log d}{N}}+\sqrt{\frac{h\log d}{n_{0}}}$. When $N\gg n_{0}$, the error rate can be significantly reduced compared to the rate obtained by only using target samples, which is in the order of $\sqrt{\frac{s\log d}{n_{0}}}+\sqrt{\frac{h\log d}{n_{0}}}$.
@@ 5.3 Minimax Lower Bounds and Optimality
$\mathcal{G}\left(s,h\right)\coloneqq\left\{\left\{\Omega^{(k)}\right\}^{K}_{k=0}:\,\Omega^{(k)}\succ 0,\,\Omega^{(k)}=\Omega^{\star}+\Gamma^{(k)\star},\,\text{supp}\left(\Omega^{\star}\right)\cap\text{supp}\left(\Gamma^{(k)\star}\right)=\emptyset\ \forall 0\leq k\leq K,\right.\\ \left.\left|\Omega^{\star}\right|_{0}\leq s,\,\max_{0\leq k\leq K}\left|\Gamma^{(k)\star}\right|_{0}\leq h,\,\max_{0\leq k\leq K}\left|\Gamma^{(k)\star}\right|_{1}\leq M_{\Gamma},\,\max_{0\leq k\leq K}\left\|\Omega^{(k)}\right\|_{2}\leq M_{\Omega},\,\max_{0\leq k\leq K}\left|\Sigma^{(k)}\right|_{\infty}\leq M_{\Sigma}\right\},$ (22)
@@
Intuitively, the performance limit of any transfer learning estimator is dictated by the information-theoretic lower bounds of estimating two parts, namely the shared component and the individual component. Hence, to derive the minimax lower bound for the transfer learning estimator, we need to provide lower bounds for estimating these two parts.
@@
The following theorem provides the minimax lower bound for estimating the shared component when all the distributions are the same.
Theorem 4.
@@
[…]
@@
The proof is based on [Cai et al., 2016b, Theorem 6.1]. See Appendix E for more details. By Theorem 4, it is easy to see that the squared Frobenius error of estimating the shared component is lower bounded by ${s\log d}/{N}$ where $N$ is the total number of samples.
@@
[…]
@@
See Appendix F for the proof, which relies on a novel construction of the packing set of the parameter space and on the celebrated Fano’s method [Wainwright, 2019, Section 15.3].
It is worth noting that Theorem 5 also provides a minimax lower bound for estimating the differential network $\Omega_{X}-\Omega_{Y}$ for two precision matrices $\Omega_{X}$ and $\Omega_{Y}$ when the $L_{1}$-norm of the differential network is bounded. To our knowledge, this is also the first lower bound for differential network estimation [Zhao et al., 2014, Yuan et al., 2017]. As a result, we can derive the first minimax optimal rate for differential network estimation. See Appendix I for a more detailed discussion.
@@
Combining Theorem 4 and Theorem 5, we have the following lower bound for any transfer learning estimator.
@@
[…]
@@
Theorem 6 demonstrates that Trans-Glasso achieves minimax optimality for the parameter space specified in (22) when $\widebar{n}\geq n_{0}$. The obtained minimax optimal rate is reasonable, considering we have $N$ samples for estimating the shared component with $s$ non-zero entries and $n_{0}$ samples for estimating the individual component with $h$ non-zero entries. Furthermore, from a practical viewpoint, the rate suggests that the target sample size only needs to be sufficiently large in relation to the sparsity level $h$ of the individual component. In contrast, if we only have target samples, the target sample size needs to be large enough to match the sparsity level $s+h$ of the entire precision matrix, which can be significantly larger.

### v2

@@ 3.2 Refinement via Differential Network Estimation
To further enhance accuracy, we refine these initial estimators by estimating the differential networks $\Psi^{(k)}=\Omega^{(k)}-\Omega^{(0)}$, which capture structural differences between each source and the target. This refinement step adjusts for potential biases in the initial estimates.
Estimating $\Psi^{(k)}$ has been extensively studied in the differential network estimation literature (50; 45), with a variety of good estimators. For instance, Yuan et al. (45) proposed estimation by solving
@@
where $\lambda^{(k)}_{\Psi}$ is a tuning parameter, and
$L_{D}\left(\Psi;\widehat{\Sigma}^{(0)},\widehat{\Sigma}^{(k)}\right)=\frac{1}{4}\left(\left\langle\widehat{\Sigma}^{(0)}\Psi,\Psi\widehat{\Sigma}^{(k)}\right\rangle+\left\langle\widehat{\Sigma}^{(k)}\Psi,\Psi\widehat{\Sigma}^{(0)}\right\rangle\right)-\left\langle\Psi,\widehat{\Sigma}^{(0)}-\widehat{\Sigma}^{(k)}\right\rangle.$
For our purpose, any reasonable differential network estimator can correct for the bias. Thus, we treat differential network estimation as a black-box algorithm, outputting $\widehat{\Psi}^{(k)}$.
@@
where $\widehat{\Psi}^{(0)}=0$ by definition. The final estimator Trans-Glasso (Transfer learning Graphical lasso) integrates both shared information and source-specific refinements, yielding a transfer learning approach that leverages structural similarities across datasets for improved precision matrix estimation. Compared to (22), our method is more sample efficient as it does not require sample splitting between the steps.
@@ 4.1 Optimization Algorithms
To implement Trans-Glasso, we first solve the Trans-MT-Glasso objective from Equation (2) to obtain initial estimates for $\Omega^{\star}$ and $\Gamma^{(k)\star}$. This is a constrained optimization problem that can be efficiently solved using the Alternating Direction Method of Multipliers (ADMM) (3). In this section, we slightly abuse notation by using superscripts to denote the iteration round and subscripts to represent the population.
@@
$X=\begin{bmatrix}\Omega_{0}\\ \Omega_{1}\\ \vdots\\ \Omega_{K}\end{bmatrix},\ Y=\begin{bmatrix}\Omega\\ \Gamma_{0}\\ \Gamma_{1}\\ \vdots\\ \Gamma_{K}\end{bmatrix},\ \text{and}\ B=\begin{bmatrix}I_{d}&I_{d}&0&\cdots&0\\ I_{d}&0&I_{d}&\cdots&0\\ \vdots&\vdots&\vdots&&\vdots\\ I_{d}&0&0&\cdots&I_{d}\end{bmatrix}$,
where $X\in\mathbb{R}^{(K+1)d\times d},Y\in\mathbb{R}^{(K+2)d\times d},B\in\mathbb{R}^{(K+1)d\times(K+2)d}$. With this notation, the objective in Equation (2) can be reformulated as
$\text{\text{minimize}}\quad f(X)+g(Y),\qquad\text{subject to}\,\,X=BY,$
@@
where $Z^{\top}=[Z^{\top}_{0},Z^{\top}_{1},\ldots,Z^{\top}_{K}]\in\mathbb{R}^{d\times(K+1)d}$ is a dual variable and $\rho$ is a penalty parameter.
@@
$Z^{(t)}=Z^{(t-1)}+\rho(X^{(t)}-BY^{(t)}).$ (10)
Note that (10) is equivalent to $Z^{(t)}_{k}=Z^{(t-1)}_{k}+\rho\left(\Omega^{(t)}_{k}-\Omega^{(t)}-\Gamma^{(t)}_{k}\right)$ for $0\leq k\leq K$. See Appendix L for detailed steps and stopping criteria.
To refine the initial estimators, we solve the D-Trace loss objective (6) for differential network estimation. Here, we use a different proximal gradient descent algorithm (31) following (47; 48).
For simplicity, let $L_{D}(\Psi)\coloneqq L_{D}\left(\Psi\,;\,\widehat{\Sigma}^{(0)},\widehat{\Sigma}^{(k)}\right)$ for the chosen $k$. In iteration $t$, we update $\Psi^{(t-1)}$ by solving
@@
where $\eta$ is a user-specified step size. Note that $\nabla L_{D}(\cdot)$ is Lipschitz continuous with constant $\|\widehat{\Sigma}^{(0)}\otimes\widehat{\Sigma}^{(k)}\|_{2}=\|\widehat{\Sigma}^{(0)}\|_{2}\|\widehat{\Sigma}^{(k)}\|_{2}$. Thus, for $0<\eta\leq\|\widehat{\Sigma}^{(0)}\|^{-1}_{2}\|\widehat{\Sigma}^{(k)}\|^{-1}_{2}$, the proximal gradient method converges (2). The update in (11) has a closed-form solution:
@@
where $A^{(t-1)}=\Psi^{(t-1)}-\eta\nabla L_{D}(\Psi^{(t-1)})$ and $x_{+}=\max\{0,x\}$ for any $x\in{\mathbb{R}}$. Details on the optimization algorithms, including stopping criteria and descriptions, are in Appendix L.
@@ 4.2 Hyperparameter Selection
following (45). After selecting $\lambda^{(k)}_{\Psi}$ and obtaining $\widehat{\Psi}^{(k)}$ for all $k\in[K]$, $\widehat{\Omega}^{(0)}$ depends on $\lambda_{\text{M}}$. Recall that $N=\sum^{K}_{k=0}n_{k}$. We choose $\lambda_{\text{M}}$ to minimize the BIC of Trans-Glasso, defined as
@@ 4.3 Identifying the Informative Set
We obtain differential network estimations $\widehat{\Psi}^{(k)}$ for all $k\in[K]$, with the hyperparameter $\lambda^{(k)}_{\Psi}$ chosen by minimizing the BIC criterion in (13). We then rank sources according to the sparsity level of $\widehat{\Psi}^{(k)}$. Let $R_{k}$ be the rank of the source $k$. For any $1\leq k_{1},k_{2}\leq K$, $|\widehat{\Psi}^{(k_{1})}|_{0}\leq|\widehat{\Psi}^{(k_{2})}|_{0}$ implies $R_{k_{1}}\leq R_{k_{2}}$. After ranking sources, we input samples into Trans-Glasso and determine the number of sources based on the cross-validation (CV) error. For $K_{\mathrm{chosen}}=0,1,\ldots,K$, we select sources $k$ with $R_{k}\leq K_{\mathrm{chosen}}$. When $K_{\mathrm{chosen}}=0$, we obtain $\widehat{\Omega}^{(0)}$ from graphical lasso (11) using the target data alone. We then compute the CV error of $K_{\mathrm{chosen}}$ by the following procedure:
@@
We set the estimated informative set as $\widehat{\mathcal{A}}=\{k\in[K]:R_{k}\leq K^{\star}_{\text{chosen}}\}$, where $K^{\star}_{\text{chosen}}=\arg\min_{k=0,1,\ldots,K}\text{CV}(k)$. Source samples from $\widehat{\mathcal{A}}$ are used to estimate $\Omega^{(0)}$. If $\widehat{\mathcal{A}}$$\emptyset$, we obtain $\widehat{\Omega}^{(0)}$ via graphical lasso on target samples. We note that sample splitting is not necessary between estimating $\mathcal{A}$ and $\Omega^{(0)}$. This procedure is called Trans-Glasso-CV. We evaluate its empirical performance in Section 6 and defer the study of its theoretical properties to future research. In the next section, we establish the theoretical results under the assumption that $\mathcal{A}=[K]$.
@@ 5 Theoretical Analysis
In this section, we establish theoretical guarantees for the Trans-Glasso algorithm. We begin by analyzing the initial estimation step using Trans-MT-Glasso, followed by an error bound for the complete Trans-Glasso estimator. Finally, we derive a minimax lower bound, demonstrating that Trans-Glasso is minimax optimal in a wide range of parameter regimes.
@@ 5.1 Analysis of Trans-MT-Glasso
The following theorem provides a high probability upper bound on the Frobenius norm error for the Trans-MT-Glasso estimator. Recall that $N=\sum^{K}_{k=0}n_{k}$ and let $\widebar{n}=N/(K+1)$.
@@
Note that the loss function defined in (3) is not strongly convex with respect to the Euclidean norm. This prevents us from obtaining error guarantee for individual precision matrix. Nevertheless, we make a key observation that the loss function exhibits strong convexity with respect to the weighted norm used in (19). See Appendix G for a detailed proof.
When we choose $\lambda_{M}\asymp\sqrt{(\log d)/N}$, the rate shown in Theorem 1 consists of two parts. The first part, which is of the order $(s\log d)/{N}$, refers to the estimation error of the shared component. In words, Trans-Glasso uses all the samples to estimate the shared component. The second part, of the order $(Kh\log d)/{N}$, relates to the estimation error of the individual components, i.e., on average, there are $N/K$ samples to estimate each individual component.
@@ 5.2 Analysis of Trans-Glasso
simultaneously for all $k=1,\ldots,K$ with probability at least $1-\delta$. We now establish a non-asymptotic error bound for this estimator, which combines the initial estimation error with the error from differential network estimation.
Corollary 1.
Let $\widehat{\Omega}^{(0)}$ be the Trans-Glasso estimator obtained in Equation (7). Under the same conditions as in Theorem 1, and assuming Equation (20) holds for the differential network estimators, with probability at least $1-2\delta$, one has
@@ 5.2.1 A differential network estimator: D-Trace loss minimization
We characterize $g^{(k)}_{\mathrm{F}}$ in the case when the D-Trace loss estimator (6) is used. We use the tighter D-Trace loss estimator analysis from (48; 47; 42) to obtain the following lemma, which is derived directly from Corollary 3 in Appendix H.
Lemma 1.
Suppose that Assumptions 1-2 hold. Assume that
$\min_{0\leq k\leq K}n_{k}\gg h^{2}\log\left(2(K+1)d^{2}/\delta\right).$
Set
@@
for the D-Trace loss estimator. Plugging the above results into Corollary 1, we have the following corollary.
@@
Let $\widehat{\Omega}^{(0)}$ be obtained by Trans-Glasso (7) with the D-Trace loss estimator used in Step 2. Assuming the assumptions in Theorem 1 and Lemma 1 hold. For a given $\delta\in(0,1]$, letting
@@
with probability at least $1-2\delta$.
The estimation error consists of three parts: shared component estimation, individual component estimation, and differential network estimation. If $\widebar{n}\geq n_{0}$ and $M_{\Gamma}$ is bounded by a universal constant, the error scales as $\sqrt{\frac{s\log d}{N}}+\sqrt{\frac{h\log d}{n_{0}}}$. Note that the rate obtained when using only target samples is on the order of $\sqrt{\frac{s\log d}{n_{0}}}+\sqrt{\frac{h\log d}{n_{0}}}$. Thus, when $s\gg h$ and $s/N\ll h/n_{0}$, transfer learning can significantly reduce the error rate. In Appendix N, we also derive an upper bound for the expected error, which follows the same order.
@@ 5.3 Minimax Lower Bounds and Optimality
$\mathcal{G}\left(s,h\right)\coloneqq\left\{\left\{\Omega^{(k)}\right\}^{K}_{k=0}:\,\Omega^{(k)}\succ 0,\,\Omega^{(k)}=\Omega^{\star}+\Gamma^{(k)\star},\,\text{supp}\left(\Omega^{\star}\right)\cap\text{supp}\left(\Gamma^{(k)\star}\right)=\emptyset\ \forall k,\right.\\ \left.\left|\Omega^{\star}\right|_{0}\leq s,\,\max_{0\leq k\leq K}\left|\Gamma^{(k)\star}\right|_{0}\leq h,\,\max_{0\leq k\leq K}\left|\Gamma^{(k)\star}\right|_{1}\leq M_{\Gamma},\,\max_{0\leq k\leq K}\left\|\Omega^{(k)}\right\|_{2}\leq M_{\Omega},\,\max_{0\leq k\leq K}\left|\Sigma^{(k)}\right|_{\infty}\leq M_{\Sigma}\right\},$ (21)
@@
Intuitively, the performance limit of estimating the target precision matrix is dictated by the information-theoretic lower bounds of estimating two parts, namely the shared component and the individual component. This motivates us to develop lower bounds for estimating these two parts.
@@
The following lemma provides the minimax lower bound for estimating the shared component when all the distributions are the same.
Lemma 2.
@@
$\Omega\in\mathcal{G}_{1}=\left\{\Omega\in\mathbb{S}^{d\times d}\,:\,\Omega\succ 0,\,\left|\Omega\right|_{0}\leq s,\,0<c_{1}\leq\gamma_{\min}(\Omega)\leq\gamma_{\max}(\Omega)\leq c_{2}<\infty\right\},$
@@
The proof is based on Theorem 6.1 in (7). See Appendix I for more details. By Lemma 2, it is easy to see that the squared Frobenius error of estimating the shared component is lower bounded by ${s\log d}/{N}$ where $N$ is the total number of samples.
@@
$\quad\quad\quad\quad\left.\left|\Delta\right|_{0}\leq h,\,\left|\Delta\right|_{1}\leq C_{\Gamma},\,0<c_{1}\leq\gamma_{\min}(\Omega)\leq\gamma_{\max}(\Omega)\leq c_{2}<\infty\right\},$
@@
See Appendix J for the proof, which relies on a novel construction of the packing set of the parameter space and on the celebrated Fano’s method (44, Section 15.3).
It is worth noting that Lemma 3 also provides a minimax lower bound for estimating the differential network $\Omega_{X}-\Omega_{Y}$ for two precision matrices $\Omega_{X}$ and $\Omega_{Y}$ when the $L_{1}$-norm of the differential network is bounded. To our knowledge, this is also the first lower bound for differential network estimation (50; 45). As a result, we can derive the first minimax optimal rate for differential network estimation. See Appendix M for a more detailed discussion.
@@
Combining Lemma 2 and Lemma 3, we have the following lower bound for estimating the target precision matrix in this transfer learning setup.
@@
$d\geq 4h,\,h\log d\geq 8\log 3,\,h\sqrt{\frac{\log d}{n_{0}}}\leq 4M_{\Gamma},\frac{h\log d}{n_{0}}\leq\min\left\{2,\,8\left(1-M_{\Omega}\right)^{2},\,8\left(1-\frac{1}{M_{\Sigma}}\right)^{2}\right\},$
@@
Theorem 2 demonstrates that Trans-Glasso achieves minimax optimality for the parameter space specified in (21) when $\widebar{n}\geq n_{0}$. The obtained minimax optimal rate is reasonable given $N$ samples for estimating the shared component with $s$ non-zero entries and $n_{0}$ samples for estimating the individual component with $h$ non-zero entries. Furthermore, from a practical viewpoint, the rate suggests that the target sample size only needs to be sufficiently large in relation to the sparsity level $h$ of the individual component. In contrast, if we only have target samples, the target sample size needs to be large enough to match the sparsity level $s+h$ of the entire precision matrix, which can be significantly larger.